Ana runs 4m, 40° North of East; 2m east; 5.2 m, 30° South of West; 6.5 m, south. Find her resultant displacement
1
Expert's answer
2020-09-15T10:04:30-0400
"A=(4\cos{40},4\sin{40})+(2,0)-\\-(5.2\cos{30},5.2\sin{30})+(0,-6.5)=\\=(0.56,-6.53)\ m"

The resultant displacement is 6.55 m, 15° East of South.